OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of -3 and contains the point (1, 1)

OpenStudy (anonymous):

y = mx +c this gives y = -3x + 4

OpenStudy (anonymous):

You can use: \[y-y_1=m(x-x_1)\] where \[\ m=-3,\ y_1=1\ and\ x_1=1\]